OpenStudy (anonymous):

For the following function, find all its critical point(s) and its absolute extrema. [ Note: If there is more than one answer enter separated by commas.] f(x)=10sqrt(x^2+10)−3x0x14

OpenStudy (anonymous):

please help because i need the answer now !

OpenStudy (anonymous):

3x0x14?

OpenStudy (anonymous):

Is that a zero-x in the second term?

OpenStudy (anonymous):

oh sorry its supposed to be 0<= x <=14

OpenStudy (anonymous):

the equation ends at ....-3x

OpenStudy (anonymous):

take the derivative. set it equal to 0. And solve for x to find the critical point

OpenStudy (anonymous):

Check this point by plugging it into the original equations. Check the boundary points 14, 0 Compare each of these and figure out which gives the biggest value.

OpenStudy (anonymous):

\[d/dx = \frac{(10 x)}{\sqrt(x^2+10)}-3 \]

OpenStudy (anonymous):

Some ugly algebra but you can soldier through it.

OpenStudy (anonymous):

can i pleaaaaase have the answer?

OpenStudy (anonymous):

x = 3sqrt(10/91) <= it's pretty ugly which makes me think you mistyped the equation